HUBBLE SPACE TELESCOPE DAILY REPORT #1543 PERIOD COVERED: 0000Z (UTC) 12/27/95 - 0000Z (UTC) 12/28/95 Daily Status Report as of 362/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets FOS/RD 6339 (Verification of the UV Spectrum of = the HST Background Light ) The Faint Object Spectrograph (Red Side) was used to make = ultraviolet observations of the HST background light. These FOS = observations were taken parallel to Hubble Deep Field (HDF) survey. Both = the HDF and these FOS data will be used to test and improve the existing = engineering model of the HST background light. The observations were = completed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 00 Successful : 00 Scheduled Reacquisitions : 07 Successful : 07 2.2 FHST Updates: Scheduled : 00 Successful : 00 2.3 Operations Notes: Temperatures are being monitored closely due to HST attitude and science = output during the Deep Field Survey. Survey ends 362:19:08z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing to report. /DEL